OpenStudy (math2400):

True or false....i believe it's false because not both intervals should be included..posting question

OpenStudy (math2400):

The series \[\sum_{n=1}^{\infty} (-2x)^{n} \] converges on the interval \[\frac{ -1 }{ 2 } < x < \frac{ 1 }{ 2 }\]

OpenStudy (zarkon):

true

OpenStudy (math2400):

it's not supposed to be 

OpenStudy (zarkon):

plug in -1/2 what do you get?

OpenStudy (math2400):

1

OpenStudy (zarkon):

\[\sum_{n=1}^{\infty} (-2(-1/2))^{n}=\sum_{n=1}^{\infty} 1=\infty\]

OpenStudy (math2400):

okkayy i see. Thank you :)